OpenStudy (anonymous):

find all the solutions of 2 cos 3x=1 (0,2pi] so i got cos 3x=1/2 or cos x=1/2 and i located it on the unit circle so i found pi/3, and 5pi/3 how do you find the other ones inbetween 0, 2pi please help

OpenStudy (anonymous):

well 2cos(3x)=1 leads to cos(3x) =1/2 as you said

OpenStudy (anonymous):

now you know that 1/2 = cos(pi/3) so you have: cos(3x) = cos(pi/3)

OpenStudy (anonymous):

so you have: 3x = pi/3 + 360k 3x = -pi/3 + 360k

OpenStudy (anonymous):

(for the general case: cos(x) = cos(a) the solutions are: x = a + 360k x = -a +360k )

OpenStudy (anonymous):

now using the equations you can get: x = pi/9 + 120k x=-pi/9 + 120k plug in here different values of k such that the x's comes in the range

OpenStudy (anonymous):

how did u get 120

OpenStudy (anonymous):

i divided both sides of the equation by 3 (so i left with x in the left side)

OpenStudy (anonymous):

look at my third response

OpenStudy (anonymous):

it was better to write 2pi/3 instead of 120 since i used pi notation im sorry for that

OpenStudy (anonymous):

ok i get that but they said the answers were( pi/9,5pi/9,7pi/9,`11pi/9,17pi/9) i got 5pi/9, pi/9 but how did they get the others

OpenStudy (anonymous):

so we have: x = pi/9 + (2pi/3)k this gives: k=0 pi/9 k=1 7pi/9 k=2 13pi/9 x=-pi/9 + (2pi/3)k k=1 5pi/9 k=2 11pi/9 k=3 17pi/9

OpenStudy (anonymous):

ok thanks